x/x+4 + 4/x+4 +2 =0

I need help on this problem.

Answer by checkley75(3666)

X/(X+4)+4/(X+4)+2=0
(X+4)/(X+4)+2=0
1+2=0
3=0 NO SOLUTION
--------------------------------------------------------
(X/X)+4+(4/X)+4+2=0
1+4+4/X+4+2=0
4/X=-11
X=-11/4